OpenStudy (anonymous):

Solving a fraction word problem using a linear equation of the form Ax=B: Charlie needs to memorize words on a vocabulary list for Latin class. He has memorized of the words, which is one-fourth of the list. How many words are on the list?

OpenStudy (anonymous):

crap that should say he memorized 12 of the words* srry typo

OpenStudy (anonymous):

Is this how I set it up....

OpenStudy (anonymous):

?

OpenStudy (anonymous):

12 words = 1/4 list

OpenStudy (anonymous):

ah nevermind I got it. it was 48 xD

OpenStudy (anonymous):

thanks anyways

OpenStudy (anonymous):

yeah he needs to learn in total 4x what he already has

OpenStudy (anonymous):

I got another one though! if u wanna help...

OpenStudy (anonymous):

just type

OpenStudy (anonymous):

The yearbook club had a meeting. The meeting had people, which is one-third of the club. How many people are in the club?

OpenStudy (anonymous):

The meeting had 21 ppl*

OpenStudy (anonymous):

21 people = 1/3 of the club

OpenStudy (anonymous):

63ppl?

OpenStudy (anonymous):

3/3 of the club =

OpenStudy (anonymous):

I think.

OpenStudy (anonymous):

yes, correct :)

OpenStudy (anonymous):

yea that's how i set it up

OpenStudy (anonymous):

omg thanks! :)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

ok one more problem

OpenStudy (anonymous):

Yoko is on her way home in her car. She has driven 24 miles so far, which is two-thirds of the way home. What is the total length of her drive?

OpenStudy (anonymous):